determine whether (-9,-3) is a solution of 8x-3y=-2
is (-9,-3) a solution to the equation?
yes or no

Answer by stanbon(75887)

determine whether (-9,-3) is a solution of 8x-3y=-2
---
Substitute x = -9 and y = -3 to get:
8*-9 - 3*-3 = -2
-72 + 9 = -2
False:

is (-9,-3) a solution to the equation?
Ans: No
==========
Cheers,
Stan H.
==========